OpenStudy (anonymous):

MEDAL TO BEST ANSWER!!!!!!!!!!!!! I need help with this equation I don't know how to solve it. Solve: 6x2 + x = 2 Please show all work

OpenStudy (johnweldon1993):

Have you worked with the Quadratic equation before?

OpenStudy (anonymous):

use the quadratic formula

OpenStudy (anonymous):

Yes but i dont know how to use it on this equation

OpenStudy (johnweldon1993):

Alright...so first we need to get your equation in the form ax^2 + bx + c = 0 How can we do that? We need to subtract 2 from both sides of your equation \[\large 6x^2 + x - 2 = 0\] Now we have a = 6 b= 1 c = -2 for your equation (these are just the coefficients of the equation)

OpenStudy (anonymous):

compare all the values of equation with standered form of qudratic equation ax+by=c

OpenStudy (anonymous):

or u can just facrosie

OpenStudy (anonymous):

6x^2+4x-3x-2=0 then take common from it 2x(3x+2)-1(3x+2)=0 then (2x-1)(3x+2)

OpenStudy (anonymous):

\[6(x +1/12)² - 49/24 = 0\]